# Double shape quantum phase transitions in the SU3-IBM (I) new $\gamma$-soft phase and the shape phase transition from the new $\gamma$-soft phase to the prolate shape

### v1

@@ Abstract
Shape quantum phase transition is an important topic in nuclear structure. In this paper, we begin to study the shape quantum phase transition in the SU3-IBM. In this new proposed model, spherical-like spectra was found to resolve the spherical nucleus puzzle, which is a new $\gamma$-soft rotational mode. In this paper, the shape phase transition along the new $\gamma$-soft line is first discussed, and then the neighbouring case at the prolate side is also studied. We find that double shape phase transitions occur along a single parameter path. The new $\gamma$-softness is really a shape phase and the shape phase transition from the new $\gamma$-soft phase to the prolate shape is found. The experimental support is also found and 108Pd is the critical nucleus.
@@ I Introduction
This simple model can describe the shape phase transitions between the spherical shape to various quadrupole deformations or among different deformed shapes (see Fig. 1 left) Casten06 ; Casten07 ; Bonatsos09 ; Casten09 ; Jolie09 ; Casten10 ; Jolos21 ; Fortunato21 ; Cejnar21 ; Jolie00 ; Cejnar03 ; Iachello04 ; Wang08 . In these studies, along a single parameter path, the shape of the nucleus changes only from one to another. After 2000, an important class of shape phase transition has attracted attentions and created controversies, which is the prolate-oblate shape phase transition Jolie01 ; Bonatsos24 . In previous IBM, the prolate-oblate shape phase transition is described via changing from the SU(3) symmetry limit to the $\overline{\textrm{SU(3)}}$ symmetry limit, and the O(6) symmetry limit is just the first-order phase transitional critical point Jolie01 , which implies that the O(6) $\gamma$-softness is not a shape phase. In this description, the spectra of the prolate and oblate shapes are the same, and it is not found in realistic nuclei. In Jolie03 , for realistic nuclei in the Hf-Hg region, the energy ratio $E_{4/2}=E_{4_{1}^{+}}/E_{2_{1}^{+}}$ of the $4_{1}^{+}$ and $2_{1}^{+}$ states is 3.33 for the prolate shape while 2.55 for the oblate shape ($E_{4/2}$ is not related to the boson number $N$). Thus this mirror symmetry appears not to exist.
Around 2020, the interacting boson model with SU(3) higher-order interactions (SU3-IBM) was proposed by one of the authors (T. Wang) Wang20 ; Wang22 , which incorporates the idea of previous IBM and the SU(3) correspondence of the rigid triaxial shape Isacker85 ; Draayer87 ; Draayer881 ; Isacker00 ; Kota20 . In this new model, the role of the SU(3) symmetry is raised to a new level, dominating all the quadrupole deformations of nuclei (see Fig. 1(a) right). It contains only the U(5) symmetry limit and the SU(3) symmetry limit. In the SU(3) symmetry limit, higher-order interactions are needed. The SU(3) second-order Casimir operator $-\hat{C}_{2}[SU(3)]$ can present the prolate shape while the SU(3) third-order Casimir operator $\hat{C}_{3}[SU(3)]$ can describe the oblate shape, which is very different from the $\overline{\textrm{SU(3)}}$ description in previous IBM. The two interactions, together with the square of the SU(3) second-order Casimir operator $\hat{C}_{2}^{2}[SU(3)]$, can describe any rigid triaxial shapes.
@@
Figure 1: (a) left represents the phase diagram of previous IBM while (a) right represents the phase diagram of the $\hat{H}$ in the SU3-IBM. The blue and red lines are two evolutional paths discussed in this paper.
@@ II Hamiltonian
here $\eta$, $\kappa$ are two controlling parameters and $c$ is the energy scale parameter. $0\leq\eta\leq 1$ and $\kappa\geq 0$. If $\eta=0$, it presents the spherical shape. If $\eta=1$ and $\kappa=0$, it describe the prolate shape. The two cases are the same as the ones in previous IBM Wang08 . If $\eta=1$ and $\kappa$ varies, this Hamiltonian describes the prolate-oblate shape phase transition Zhang12 , which is a finite $N$ effect.
Fig. 1(a) left shows the phase diagram of previous IBM in the large-$N$. Above the real line, the spherical shape exists, and under the real line, the deformed shapes exist. The deformed region is divided by the blue line which is part of the connected line between the U(5) symmetry limit to the O(6) symmetry limit. The left part of the blue line presents the prolate shape while the right part presents the oblate shape. The blue line is the critical line between the prolate and oblate shapes. The crossover point of the green line and the blue line is the triple point with the spherical, prolate and oblate shapes Casten02 ; Warner02 . Obviously, the O(6) $\gamma$-softness in not a shape phase.
Figure 2: The evolutional behaviors of the ground state in $\hat{H}$ for
@@
In this paper, the shape phase transition along the real blue line is studied, and not stress the degenerate line, which is difficult to discuss. The SU(3) degenerate point is at $\kappa_{0}=\frac{3N}{2N+3}$ and the red point is at $0.9\kappa_{0}$. Through previous analysis, the red point presents the prolate shape. This is very interesting. For the real red line from the U(5) symmetry limit to the red point, intuitively, the shape changes from the spherical shape to the prolate. Through later numerical calculations, we find that this shape transition is not direct but through the new $\gamma$-softness region, and the double shape quantum phase transition can occur. The right part of the real blue line is not discussed in this paper, and will be studied in next paper for investigating the scope of the new $\gamma$-soft region.
It is important to emphasize here why the O(6) $\gamma$-softness in previous IBM and actual $\gamma$-soft nuclei do not match. In actual nuclei, there exists many nuclei in the $\gamma$-soft region, such as Pt, Xe nuclei, so it is hard to believe that this is just the O(6)-softness in previous IBM, a critical point of shape phase transition. In the SU3-IBM, such a conceptual conflict does not exist. We believe that the shape phase transition given by the SU3-IBM is an accurate description of the realistic shape phase transitions in nuclei. This point has been preliminarily confirmed by the prolate-oblate asymmetric shape phase transition in the Hf-Hg region Wang23 .
@@ III Double shape quantum phase transition
Figure 3: (a) The evolutional behaviors of the partial low-lying levels as a function of $\chi$ for $N=9$ from the SU(3) symmetry limit to the O(6) symmetry limit; (b) The evolutional behaviors of the $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$ (blue line) , $B(E2;4_{1}^{+}\rightarrow 2_{1}^{+})$ (red line), $B(E2;6_{1}^{+}\rightarrow 4_{1}^{+})$ (green line) as a function of $\chi$. The parameters are deduced from Pan24 .
Figure 4: (a) The evolutional behaviors of the partial low-lying levels as a function of $\chi$ for $N=9$ from the SU(3) symmetry limit to the O(6) symmetry limit; (b) The evolutional behaviors of the $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$ (blue line) , $B(E2;4_{1}^{+}\rightarrow 2_{1}^{+})$ (red line), $B(E2;6_{1}^{+}\rightarrow 4_{1}^{+})$ (green line) as a function of $\chi$. The parameters are deduced from Pan24 .
@@
Fig. 2(a) also shows the energy evolution of the ground state of $\hat{H}$ along the real red line in Fig. 1(b) for $N=10$ (dashed red line) and for $N=60$ (real red line). For $N=60$, it is shown that, around $\eta=0.5$, a new phase transition point appears. The part of the real red line deviating from the real blue line is prominent, which has a steeper descent. Between $\eta=0$ and $\eta=0.2$, there is the spherical shape, and between $\eta=0.5$ and $\eta=1$, there is the prolate shape. Obviously, between the two shapes, the new $\gamma$-softness exists. When $\kappa$ changes from $\kappa_{0}$ to $0.9\kappa_{0}$, the new $\gamma$-soft region reduces, but it does exist. A key point is that, between $\eta=0.2$ and $\eta=0.5$, the red and blue lines are nearly degenerate, so when $\kappa$ changes from $\kappa_{0}$ to $0.9\kappa_{0}$, the energies of the new $\gamma$-softness are nearly the same and not reduce. This implies that, the new $\gamma$-softness is really a phase shape. Thus along the left real red line, double shape quantum phase transition can occur. This can not occur for the shape phase transition along the real blue line and in previous IBM.
Figure 5: (a) The evolutional behaviors of the partial low-lying levels as a function of $\chi$ for $N=9$ from the SU(3) symmetry limit to the O(6) symmetry limit; (b) The evolutional behaviors of the $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$ (blue line) , $B(E2;4_{1}^{+}\rightarrow 2_{1}^{+})$ (red line), $B(E2;6_{1}^{+}\rightarrow 4_{1}^{+})$ (green line) as a function of $\chi$. The parameters are deduced from Pan24 .
Figure 6: (a) The evolutional behaviors of the partial low-lying levels as a function of $\chi$ for $N=9$ from the SU(3) symmetry limit to the O(6) symmetry limit; (b) The evolutional behaviors of the $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$ (blue line) , $B(E2;4_{1}^{+}\rightarrow 2_{1}^{+})$ (red line), $B(E2;6_{1}^{+}\rightarrow 4_{1}^{+})$ (green line) as a function of $\chi$. The parameters are deduced from Pan24 .
In previous IBM, similar result of the ground energy evolution along the real blue line can be also obtained from the U(5) symmetry limit to the O(6) symmetry limit. If the O(6) symmetry limit is the prolate-oblate critical point, the connected line in the deformed region between the U(5) symmetry limit and the O(6) symmetry limit is a prolate-oblate critical line. When deviating from the critical line, the energies of the deformed region reduce and the double shape phase transitions can not be observed Wang08 .
@@
Figure 7: (a) The evolutional behaviors of the partial low-lying levels as a function of $\chi$ for $N=9$ from the SU(3) symmetry limit to the O(6) symmetry limit; (b) The evolutional behaviors of the $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$ (blue line) , $B(E2;4_{1}^{+}\rightarrow 2_{1}^{+})$ (red line), $B(E2;6_{1}^{+}\rightarrow 4_{1}^{+})$ (green line) as a function of $\chi$. The parameters are deduced from Pan24 .
@@
Fig. 4(a) presents the evolutional behaviors of the partial low-lying states for $N=35$ along the real blue line in Fig. 1(b), which is a new result. The shape phase transition from the spherical shape to the new $\gamma$-soft rotation becomes more prominent. When $\eta>2$, the $4_{1}^{+}$ and $2_{2}^{+}$ states begin to separate because the degenerate line curves to the left. Besides, the level-anticrossing of the $0_{2}^{+}$, $0_{3}^{+}$, $0_{4}^{+}$ states becomes more clear. Fig. 4(b) shows the evolutional behaviors of the B(E2) values of the $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$, $B(E2;4_{1}^{+}\rightarrow 2_{1}^{+})$, $B(E2;2_{2}^{+}\rightarrow 2_{1}^{+})$, $B(E2;0_{2}^{+}\rightarrow 2_{1}^{+})$ along the real blue line in Fig. 1(b). The shape phase transition becomes more clear.
Now we discuss the double shape phase transitions along the real red line in Fig. 1(b). Fig. 5(a) presents the evolutional behaviors of the partial low-lying states for $N=35$. Between the $\eta=0.2$ and $\eta=0.5$, the $4_{1}^{+}$ and $2_{2}^{+}$ are nearly degenerate, and the spectra are also similar to the spherical-like spectra, so this is the new $\gamma$-soft phase. When $\eta>0.5$, obviously it is the prolate shape. In Fig. 6(a) the case of $N=35$ is shown and the shape phase transition from the $\gamma$-soft rotation to the prolate shape becomes very prominent. Thus $k=0.9\kappa_{0}$ and $\eta=0.5$ is the phase transition critical point.
@@
Finally, we look for some experimental evidences for the existence of the shape phase transition from the new $\gamma$-soft phase to the prolate shape. Fig. 8(a) shows the different evolutional trends of the $Q_{2_{1}^{+}}$ values of 108-114Cd and 104-110Pd. This was first observed in Wang25 , and can be regarded as the first strong support for the existence of the spherical-like spectra. The discussions in this paper further support this conclusion and is an indirect experimental support for the existence of the shape phase transition. A direct support can be observed in Fig. 8(b). In the previous analysis, we see that, from the new $\gamma$-soft phase to the prolate shape, the value of $B(E2;0_{2}^{+}\rightarrow 2_{1}^{+})$ increases first, and then decreases. here we define $\beta=B(E2;0_{2}^{+}\rightarrow 2_{1}^{+})/B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$. Fig. 8(b) presents the $\beta$ evolution of 104-110Pd and obviously it is clearly in line with the theoretical prediction. Thus 104,106Pd are two typical new $\gamma$-soft nuclei and 108Pd is a critical nucleus from the new $\gamma$-soft phase to the prolate shape. However for finite-$N$, the spherical-like spectra and the critical spectra may be not distinguished. A detailed discussions on the properties of 104-110Pd will be given in a future paper.
It should be noticed that this double shape quantum phase transitions cannot be verified directly. In the last two decades, the experimental discovery that nuclei previously thought to be spherical cannot be confirmed to be spherical is a breakthrough in the field of nuclear structure. If the spherical nucleus is absent, it is difficult to confirm the shape phase transition from the spherical shape to the new $\gamma$-soft phase. In our discussion, we also found that the spherical nucleus and the critical nucleus at $\eta=0.2$ are also difficult to be distinguished, which can help us further discuss those nuclei that look like spherical.

### v2

# Double shape quantum phase transitions in the SU3-IBM: new $\gamma$-soft phase and the shape phase transition from the new $\gamma$-soft phase to the prolate shape

@@ Abstract
Shape quantum phase transition is an important topic in nuclear structure. In this paper, we begin to study the shape quantum phase transition in the SU3-IBM. In this new proposed model, spherical-like spectra was found to resolve the spherical nucleus puzzle, which is a new $\gamma$-soft rotational mode. In this paper, the shape phase transition along the new $\gamma$-soft line is first discussed, and then the neighbouring case at the prolate side is also studied. We find that double shape phase transitions occur along a single parameter path. The new $\gamma$-softness is really a shape phase and the shape phase transition from the new $\gamma$-soft phase to the prolate shape is found. The experimental support is also found and 108Pd may be the critical nucleus.
@@ I Introduction
This simple model can describe the shape phase transitions between the spherical shape to various quadrupole deformations or among different deformed shapes (see Fig. 1(a) left) Casten06 ; Casten07 ; Bonatsos09 ; Casten09 ; Jolie09 ; Casten10 ; Jolos21 ; Fortunato21 ; Cejnar21 ; Jolie00 ; Cejnar03 ; Iachello04 ; Wang08 . In these studies, along a single parameter path, the shape of the nucleus changes only from one to another. After 2000, an important class of shape phase transition has attracted attentions and created controversies, which is the prolate-oblate shape phase transition Jolie01 ; Bonatsos24 . In previous IBM, the prolate-oblate shape phase transition is described via changing from the SU(3) symmetry limit to the $\overline{\textrm{SU(3)}}$ symmetry limit, and the O(6) symmetry limit is just the first-order phase transitional critical point Jolie01 , which implies that the O(6) $\gamma$-softness is not a shape phase. In this description, the spectra of the prolate and oblate shapes are the same, and it is not found in realistic nuclei. In Jolie03 , for realistic nuclei in the Hf-Hg region, the energy ratio $E_{4/2}=E_{4_{1}^{+}}/E_{2_{1}^{+}}$ of the $4_{1}^{+}$ and $2_{1}^{+}$ states is 3.33 for the prolate shape while 2.55 for the oblate shape ($E_{4/2}$ is not related to the boson number $N$). Thus this mirror symmetry appears not to exist.
Around 2020, an extension of the interacting boson model with SU(3) higher-order interactions (SU3-IBM) was proposed by one of the authors (T. Wang) Wang20 ; Wang22 , which incorporates the idea of previous IBM and the SU(3) correspondence of the rigid triaxial shape Isacker85 ; Draayer87 ; Draayer881 ; Isacker00 ; Kota20 . In this new model, the role of the SU(3) symmetry is raised to a new level, dominating all the quadrupole deformations of nuclei (see Fig. 1(a) right). It contains only the U(5) symmetry limit and the SU(3) symmetry limit. In the SU(3) symmetry limit, higher-order interactions are needed. The SU(3) second-order Casimir operator $-\hat{C}_{2}[SU(3)]$ can present the prolate shape while the SU(3) third-order Casimir operator $\hat{C}_{3}[SU(3)]$ can describe the oblate shape, which is very different from the $\overline{\textrm{SU(3)}}$ description in previous IBM. The two interactions, together with the square of the SU(3) second-order Casimir operator $\hat{C}_{2}^{2}[SU(3)]$, can describe any rigid triaxial shapes.
@@
Figure 1: (a) left represents the phase diagram of previous IBM while (a) right represents the phase diagram of the $\hat{H}$ in the SU3-IBM. In (b) the real blue and real red lines are two evolutional paths discussed in this paper. Along the real red line, the double shape phase transitions can occur.
@@ II Hamiltonian
here $\eta$, $\kappa$ are two controlling parameters and $c$ is the energy scale parameter. $0\leq\eta\leq 1$ and $\kappa\geq 0$. If $\eta=0$, it presents the spherical shape. If $\eta=1$ and $\kappa=0$, it describe the prolate shape. The two cases are the same as the ones in previous IBM Wang08 . If $\eta=1$ and $\kappa$ varies, this Hamiltonian describes the prolate-oblate shape phase transition Zhang12 , which is a finite-$N$ effect.
Fig. 1(a) left shows the phase diagram of previous IBM in the large-$N$ limit. Above the real green line, the spherical shape exists, and under the real green line, the deformed shapes exist. The deformed region is divided by the real blue line which is part of the connected line between the U(5) symmetry limit to the O(6) symmetry limit. The left part of the blue line presents the prolate shape while the right part presents the oblate shape. The blue line is the critical line between the prolate and oblate shapes. The crossover point of the green line and the blue line is the triple point with the spherical, prolate and oblate shapes Casten02 ; Warner02 . Obviously, the O(6) $\gamma$-softness in not a shape phase. Up to two-body interactions, the IBM can not describe the rigid triaxial shapes Chen1981 .
Figure 2: (a) The energy evolution of the ground state of $\hat{H}$ along the real blue (red) line in Fig. 1(b) for $N=10$ (dashed blue (red) line) and for $N=60$ (real blue (red) line). (a) The $\bar{n}_{d}$ evolution of the ground state of $\hat{H}$ along the real blue (red) line in Fig. 1(b) for $N=10$ (dashed blue (red) line) and for $N=60$ (real blue (red) line).
@@
In this paper, the shape phase transition along the real blue line is studied, and not stress the degenerate line, which is difficult to discuss. The SU(3) degenerate point is at $\kappa_{0}=\frac{3N}{2N+3}$ and the red point is at $0.9\kappa_{0}$. Through previous analysis, the red point presents the prolate shape. This is very interesting. For the real red line from the U(5) symmetry limit to the red point, intuitively, the shape changes from the spherical shape to the prolate. Through later numerical calculations, we find that this shape transition is not direct but through the new $\gamma$-softness region, and the double shape quantum phase transition can occur. The right part of the real blue line is not discussed in this paper, and will be studied in next paper for investigating the scope of the new $\gamma$-soft region Zhao2 .
It is important to emphasize here why the O(6) $\gamma$-softness in previous IBM and actual $\gamma$-soft nuclei do not match. In actual nuclei, there exists many nuclei in the $\gamma$-soft region, such as Os, Pt, Xe, Ba nuclei ensdf , so it is hard to believe that this is just the O(6)-softness in previous IBM, a critical point of shape phase transition. In the SU3-IBM, such a conceptual conflict does not exist. We believe that the shape phase transition given by the SU3-IBM is an accurate description of the realistic shape phase transitions in nuclei. This point has been preliminarily confirmed by the prolate-oblate asymmetric shape phase transition in the Hf-Hg region Wang23 .
@@
Figure 3: (a) The evolutional behaviors of the partial low-lying levels as a function of $\eta$ for $N=10$ along the real blue line in Fig. 1(b); (b) The evolutional behaviors of the $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$ (black line), $B(E2;4_{1}^{+}\rightarrow 2_{1}^{+})$ (blue line), $B(E2;2_{2}^{+}\rightarrow 2_{1}^{+})$ (red line), $B(E2;0_{2}^{+}\rightarrow 2_{1}^{+})$ (green line) along the real blue line in Fig. 1(b).
@@ III Double shape quantum phase transition
Figure 4: (a) The evolutional behaviors of the partial low-lying levels as a function of $\eta$ for $N=35$ along the real blue line in Fig. 1(b); (b) The evolutional behaviors of the $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$ (black line), $B(E2;4_{1}^{+}\rightarrow 2_{1}^{+})$ (blue line), $B(E2;2_{2}^{+}\rightarrow 2_{1}^{+})$ (red line), $B(E2;0_{2}^{+}\rightarrow 2_{1}^{+})$ (green line) along the real blue line in Fig. 1(b).
@@
Figure 5: (a) The evolutional behaviors of the partial low-lying levels as a function of $\eta$ for $N=10$ along the real red line in Fig. 1(b); (b) The evolutional behaviors of the $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$ (black line), $B(E2;4_{1}^{+}\rightarrow 2_{1}^{+})$ (blue line), $B(E2;2_{2}^{+}\rightarrow 2_{1}^{+})$ (red line), $B(E2;0_{2}^{+}\rightarrow 2_{1}^{+})$ (green line) along the real red line in Fig. 1(b).
Fig. 2(a) also shows the energy evolution of the ground state of $\hat{H}$ along the real red line in Fig. 1(b) for $N=10$ (dashed red line) and for $N=60$ (real red line). For $N=60$, it is shown that, around $\eta=0.5$ (denoted by the middle dashed line), a new shape phase transition appears. The part of the real red line deviating from the real blue line is prominent, which has a steeper descent. Between $\eta=0$ and $\eta=0.2$, there is the spherical shape, and between $\eta=0.5$ and $\eta=1$, there is the prolate shape. Obviously, between the two shapes, the new $\gamma$-softness exists. When $\kappa$ changes from $\kappa_{0}$ to $0.9\kappa_{0}$, the new $\gamma$-soft region reduces, but it does exist. A key point is that, between $\eta=0.2$ and $\eta=0.5$, the red and blue lines are nearly degenerate, so when $\kappa$ changes from $\kappa_{0}$ to $0.9\kappa_{0}$, the energies of the new $\gamma$-softness are nearly the same and do not reduce. This implies that, the new $\gamma$-softness is really a phase shape. Thus along the left real red line, double shape quantum phase transition can occur. This can not occur for the shape phase transition along the real blue line and in previous IBM.
Some details need further elaboration. For $N=10$, the real blue line in Fig. 1(b) is nearly the critical line between the prolate shape and the new $\gamma$-soft region, so at the left side of the real blue line, it seems only the prolate shape region. This implies that the new $\gamma$-soft region is larger than expected. which will be further clarified in Zhao2 . For $N=35$, the critical line curves to the left, so the new shape phase transition becomes more prominent.
Figure 6: (a) The evolutional behaviors of the partial low-lying levels as a function of $\eta$ for $N=35$ along the real red line in Fig. 1(b); (b) The evolutional behaviors of the $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$ (black line), $B(E2;4_{1}^{+}\rightarrow 2_{1}^{+})$ (blue line), $B(E2;2_{2}^{+}\rightarrow 2_{1}^{+})$ (red line), $B(E2;0_{2}^{+}\rightarrow 2_{1}^{+})$ (green line) along the real red line in Fig. 1(b).
In previous IBM, similar result of the ground energy evolution along the real blue line can be also obtained along the evolution path from the U(5) symmetry limit to the O(6) symmetry limit. If the O(6) symmetry limit is the prolate-oblate critical point, the connected line in the deformed region between the U(5) symmetry limit and the O(6) symmetry limit is a prolate-oblate critical line. When deviating from the critical line, the energies of the deformed region reduce and the double shape phase transitions can not be observed Wang08 .
Figure 7: The $Q_{2_{1}^{+}}$ evolution of the $2_{1}^{+}$ state of $\hat{H}$ along the real blue (red) line in Fig. 1(b) for $N=7$ (dotted blue (red) line), $N=10$ (dashed blue (red) line) and for $N=35$ (real blue (red) line).
@@
Fig. 4(a) presents the evolutional behaviors of the partial low-lying states for $N=35$ along the real blue line in Fig. 1(b), which is a new result. The shape phase transition from the spherical shape to the new $\gamma$-soft rotation becomes more prominent. When $\eta>0.2$, the $4_{1}^{+}$ and $2_{2}^{+}$ states begin to separate because the degenerate line curves to the left. Besides, the level-anticrossing of the $0_{2}^{+}$, $0_{3}^{+}$, $0_{4}^{+}$ states becomes more clear Li25 ; Jolie02 ; Arias03 . Fig. 4(b) shows the evolutional behaviors of the B(E2) values of the $B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$, $B(E2;4_{1}^{+}\rightarrow 2_{1}^{+})$, $B(E2;2_{2}^{+}\rightarrow 2_{1}^{+})$, $B(E2;0_{2}^{+}\rightarrow 2_{1}^{+})$ along the real blue line in Fig. 1(b). The shape phase transition becomes more clear. When $N$ increases, the values of $B(E2;2_{2}^{+}\rightarrow 2_{1}^{+})$, $B(E2;0_{2}^{+}\rightarrow 2_{1}^{+})$ becomes smaller.
Now we discuss the double shape phase transitions along the real red line in Fig. 1(b). Fig. 5(a) presents the evolutional behaviors of the partial low-lying states for $N=10$. Between the $\eta=0.2$ and $\eta=0.5$, the $4_{1}^{+}$ and $2_{2}^{+}$ are nearly degenerate, and the spectra are also similar to the spherical-like spectra, so this is the new $\gamma$-soft phase. When $\eta>0.5$, obviously it is the prolate shape. In Fig. 6(a) the case of $N=35$ is shown and the shape phase transition from the $\gamma$-soft rotation to the prolate shape becomes very prominent. Thus $k=0.9\kappa_{0}$ and $\eta=0.5$ is the phase transition critical point.
@@
Finally, we look for some experimental evidences for the existence of the shape phase transition from the new $\gamma$-soft phase to the prolate shape. Fig. 8(a) shows the different evolutional trends of the $Q_{2_{1}^{+}}$ values of 108-114Cd and 104-110Pd. This was first observed in Wang25 , and can be regarded as the first strong support for the existence of the spherical-like spectra. The discussions in this paper further support this conclusion and is an indirect experimental support for the existence of the shape phase transition. A direct support can be observed in Fig. 8(b). In the previous analysis, we see that, from the new $\gamma$-soft phase to the prolate shape, the value of $B(E2;0_{2}^{+}\rightarrow 2_{1}^{+})$ increases first, and then decreases. here we define $\beta=B(E2;0_{2}^{+}\rightarrow 2_{1}^{+})/B(E2;2_{1}^{+}\rightarrow 0_{1}^{+})$. Fig. 8(b) presents the $\beta$ evolution of 104-110Pd and obviously it is clearly in line with the theoretical prediction. Thus 104,106Pd are two typical new $\gamma$-soft nuclei WangPd and 108Pd may be a critical nucleus from the new $\gamma$-soft phase to the prolate shape. However for finite-$N$, the spherical-like spectra and the critical spectra may be not distinguished. A detailed discussions on the properties of 104-110Pd will be given in a future paper.
It should be noticed that this double shape quantum phase transitions cannot be verified directly. In the last two decades, the experimental discovery that nuclei previously thought to be spherical cannot be confirmed to be spherical is a breakthrough in the field of nuclear structure Garrett18 . If the spherical nucleus is absent, it is difficult to confirm the shape phase transition from the spherical shape to the new $\gamma$-soft phase. In our discussion, we also found that the spherical nucleus and the critical nucleus at $\eta=0.2$ are also difficult to be distinguished, which can help us further discuss those nuclei that look like spherical.